OpenStudy (anonymous):

@ ..if 12.5 cu reacts with excess cl gas then 25.4g of cu(ii)cl are produced..find the theoretical and percentage yield

OpenStudy (anonymous):

cu + cl2 = cucl2 63.55 + 2* 35.45 = 134.45 - taking relative atomic masses 12.5 g theoretically produces ( 134.45 / 63.55 ) * 12.5 = 26.446 g persentage yield = (25.4 /26.446) * 100 = 96.1 % to 1 dec place

OpenStudy (preetha):

Good answer.